INAF LBC+MODS queue observing, 2012 February 23-24

Observer: K. Boutsia, M. Fumana
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LBC+MODS

Summary

The sky is covered with thin cirrus before sunset. No use taking flats.
Starting with LBC to try again the NS target. Today the seeing was better and we observed both pointing (East and West).
The target IC10VI is already at very low elevation just after twilight and to have it within specs we need exceptional seeing. No such luck tonight.
At 04:00UT we switched to MODS and we completed 4 projects: Nonino (macs0647), Fiore mask1 and mask2, Donnarumma (Abell mask1).
We also started Donnarumma (Abell mask2). We did 3 out of 4 exposures for this mask. The 3rd exposure finished within the twilight. At that time we were pointing NW. The sky in the spectra is indeed higher than the previous ones.
The seeing has been variable but during the MODS time it has been mostly <1.3". For the seeing during each target see figures at the end of the log.
The calibrations for the MODS data are complete (flats + lamps and standard star for all masks).
The Donnarumma calibrations for mask1 and mask2 have as PROPID sdss1029 even though the actual target is A1703. The masks are the correct ones!!

Data Summary

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan 1hr seeing was better than yesterday and we did both pointings
Storm IC10VI V-bess / I-bess 5min seeing was ~2.0", measured on the images, so we aborted
Nonino macsj0647 dual grating 1hr seeing was ~1.3 which is comparable to the rest of the 4hrs - this mask is considered complete
Fiore Cosmos1 dual grating 0.5hr seeing was <1.3" - this mask is considered complete
Fiore Cosmos2 dual grating 1hr 0.5hr was within specs and the second 0.5hr was as yesterday - this mask is considered complete
Donnarumma A1703m1 red grating 1hr50min seeing was around 1.0" for most of the time - this mask is considered complete
Donnarumma A1703m2 red grating 1hr50min seeing was <1.3" - the last exposure finished after sky started getting brighter - sky in the spectra does look higher!!
STD_MODS Feige34 dual/red grating good for all MODS programs

Details

01:55 Slew to 20120223F
  • dofpia,/First,/X2
4251      0   -116  -1624     18    220     80   1.56   1.83     red
8239      0      0      0      0      0      0   2.68   3.25    blue

-503     98     61    -26    395    278     95   1.87   0.33     red
 -1544   -357   -324   -679    -95    340    100   2.33   0.73    blue

 209   -105    114   -123   -130    196     62   1.76   0.19     red
 -343   -112   -270   -383   -114    335     98   2.39   0.33    blue
  • lbcrangebal
  • dimm seeing 1.9" -2.0"

02:07 20120224W - take one field image with normal guiding to see what the seeing is like.
  • seeing measured in the image 1.4"-1.6"
  • dimm seeing 1.4"
  • dofpia again
 1233    104     40    117     31   -184    -39   1.23   0.51     red
 1866   -174    -45    -58     -8    -45      6   1.82   0.74    blue

  442     35     85   -177   -133   -112    -26   1.01   0.23     red
 412     56    105   -113     -3   -103     -5   1.74   0.20    blue

 -319    -60     63    110     48   -125    -14   1.10   0.18     red
358   -125    -64    114    -22    -83    -24   1.65   0.18    blue

02:21 Set NSIGUI and then send 20120224W without guiding

03:04 slewing to the IC10VI
  • if we wait until the NEO is done it will be too low in elevation
  • dofpia
 -6376      0      0      0      0      0      0   1.40   2.51     red
-5823    -21   -172     53    320    249     62   1.57   2.31    blue

 848    -51    164    282   -223     96      2   1.46   0.39     red
  -62   -131     13     29     46     70     -4   1.52   0.10    blue

-83     96    134    109     49     96     30   1.62   0.13     red
 -481     39     79     -5     57    128     28   1.70   0.22    blue

03:11 IC10VI
  • the red filter got stuck - we had to stop the OB and resend it
  • the seeing on the blue images was 1.8" - on the red also
  • we tried - but data is out of spec - we abort and move back to the NSI target

03:20 Slew back to 1999AO10 for the east field
  • dofpia
1075     38   -144    258    -22     18     44   1.86   0.45     red
2616   -163    107     24     70     58     39   1.93   1.04    blue

 -102     19   -120   -188     50     42     29   1.58   0.13     red
 436    -98     47     83    -18    -53      5   1.63   0.20    blue

03:27 20120224E
  • dimm seeing 1.7"

04:09 SWITCH TO MODS

04:30 Pointing correction and collimation

04:49 acqMODS macsj0647_mods.acq mods1r 0005-0007
   dX:  -0.008 arcsec
   dY:  -0.846 arcsec
  dPA: -0.1204 degrees

05:02 execMODS macsj0647_mods.obs mods1b 0005-0008 / mods1r 0008-0011
  • seeing on GCS 1.1"-1.4" - this mask is considered completed!!
  • in the header there is not the MASKNAME!!!!

06:11 acqMODS mods.56.FIORE1.acq mods1r 0012-0014
   dX:  -0.100 arcsec
   dY:  -0.708 arcsec
  dPA: -0.1016 degrees

06:25 execMODS mods.56.FIORE1.obs mods1b 0009 / mods1r 0015
  • dimm seeing 1.2"

07:00 acqMODS mods.57.FIORE2.acq mods1r 0016-0018
   dX:   0.122 arcsec
   dY:  -1.125 arcsec
  dPA: -0.1832 degrees

07:11 execMODS mods.57.FIORE2.obs mods1b 0010-0011 / mods1r 0019-0020
  • the first 0.5hr is within specs but when the second has started the seeing went up to 1.8"
  • dimm seeing 2.0' for the second exposure
  • humidity is going up - cirrus passing?

08:16 Abort because the seeing is worse than yesterday - not worth accumulating hrs.
  • it has never started obtaining the second image for the blue side! - actually it did - it is the gui that did not refresh!!
  • pc is hang - after aborting the blue side is not responding! - blue reset in the command line - and press the refresh button in the gui.

08:20 acqMODS Feige34_std.acq mods1r 0021-0026
  • offsetxy 0.149 11.023 rel
  • we still loose the guide star every now and then, probably there is cirrus passing by.
  • changing the exptime on the guider from 2sec to 4sec

08:34 execMODS Feige34_std.obs mods1b 0012 / mods1r 0027
  • this is with dual_grating - suitable for Nonino and Fiore

08:42 execMODS Feige34_stdRED.obs mods1r 0028
  • this is with dual_grating - suitable for Donnarumma

08:47 acqMODS mods.201151.newmi1e.acq mods1r 0029-0031
   dX:  -0.115 arcsec
   dY:  -0.849 arcsec
  dPA: -0.1092 degrees
  • measured seeing on the acquisition image is 1.7"

09:01 execMODS mods.201151.newmi1e.obs mods1r 0032-0034
  • Completed after 3 exposures (2250s each)
  • the seeing improved a lot after the first 15min and it stayed around 1.0"
  • that's why we did one more exposure so that adding the data of previous nights we have 2hr30min within specs.

11:01 Slew to mods.201150.newmi2e.acq mods1r 0035-0037
  • acqMODS mods.201150.newmi2e.acq
   dX:  -0.232 arcsec
   dY:  -0.992 arcsec
  dPA: -0.0859 degrees

11:12 execMODS mods.201150.newmi2e.obs mods1r 0038-0040

13:11 Abort the script after the 3rd exposure
  • the last exposure mods1r 0040 finishes 20min after the sky started getting brighter
  • we are facing NW which should be the darkest part of the sky at this hour.

13:14 Close Dome - Telescope to zenith for calibrations
  • The Donnarumma calibrations for mask2 has as PROPID sdss1029 even though the actual target is A1703. The maskid is the correct one!!
  • execMODS grmoslampsDonnaRm2.cal mods1r 0041-0043
  • execMODS grmosflatsDonnaRm2.cal mods1r 0044-0053

13:51 END OF NIGHT
  • execMODS ~MODSeng/modsScripts/Support/modsSleep.pro

  • Seeing reported by GCS during Nonino 0647:
    Nonino0224.jpg

  • Seeing reported by GCS during Fiore1:
    Fiore0224.jpg

  • Seeing reported by GCS during Fiore2:
    Fiorem2UT0224.jpg

  • Seeing reported by GCS during Donnarumma M1:
    DonM10224.jpg

  • Seeing reported by GCS during Donnarumma M2:
    DonM2UT0224.jpg

-- KonstantinaBoutsia - 24 Feb 2012
Topic revision: r5 - 24 Feb 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback